if P(E) = 0.40, P(E or F) = 0.50, and P(E and F) = 0.05, find P(F)
Answer by ikleyn(52790)

.
if P(E) = 0.40, P(E or F) = 0.50, and P(E and F) = 0.05, find P(F)
~~~~~~~~~~~~~~

Use the general formula of the probability theory

    P(E or F) = P(E) + P(F) - P(E and F).


Substitute the given values into the formula.  You will get this equation

    0.50 = 0.40 + P(F) - 0.05.


In this equation, only value  P(F) is unknown.  Find it from the equation

    P(F) = 0.50 - 0.40 + 0.05 = 0.15.    ANSWER

Solved.